f(x)=x^2-3 the domain is {-1/2,1/4,1/2}. Determine the range. Thanks for your help.
Answer by Edwin McCravy(20054)

f(x)=x^2-3 the domain is {-1/2,1/4,1/2}

domain                                                             range
value                                                              value 
  x     f(x)   =    x²   - 3 =       simplification             =  f(x)
-----------------------------------------------------------------------
-1/2   f(-1/2) = (-1/2)² - 3 = 1/4  - 3 = 3/12 - 36/12 = -33/12 = -11/4
 1/4   f(1/4)  = (1/4)²  - 3 = 1/16 - 3 = 1/16 - 48/16 =         -47/16
 1/2   f(1/2)  = (1/2)²  - 3 = 1/4  - 3 = 3/12 - 36/12 = -33/12 = -11/4 

The range is {-11/4, -47/16}  [we do not list -11/4 but once.]

Edwin